OpenStudy (okdutchman7):

I need help. WILL MEDAL! @ganeshie8 @Hero @TheSmartOne @Preetha @zepdrix @mathmale @rebeccaxhawaii @Vocaloid @dtan5457 @MrNood @amoodarya

OpenStudy (mathmale):

How's that again? sort3?

OpenStudy (okdutchman7):

\[-2 + \sqrt{3} \sec \theta = 0\]

OpenStudy (mathmale):

\[-2 + \sqrt{3} \sec \theta = 0\]

OpenStudy (mathmale):

First of all, solve for sec theta. Your result?

OpenStudy (okdutchman7):

that's right

OpenStudy (okdutchman7):

im kinda confused on this type of problem. i dont know where to start

OpenStudy (mathmale):

How would you solve for x here?\[-2 + \sqrt{3} x = 0\]

OpenStudy (okdutchman7):

\[Move \sqrt{3} x \to the other side\]